OpenStudy (anonymous):

Simplify [sin(s-t)/sin t] + [cos (s-t)/cos t] to sin s/ (sin s cos t) I used the identity of both but I can't seem to get rid of any of the functions

myininaya (myininaya):

\[\frac{\sin(s-t)}{sint}+\frac{\cos(s-t)}{cost}=\frac{sins*cost-sint*coss}{sint}+\frac{coss*cost+sins*sint}{cost}\] =\[\frac{sins*cost}{sint}-\frac{sint*cost}{sint}+\frac{coss*cost}{cost}+\frac{sins*sint}{cost}\] =\[\frac{sins*cost}{sint}-cost+cost+\frac{sins*sint}{cost}\] =\[\frac{cost(sins*cost)+sint(sins*sint)}{sint*cost}=\frac{sins(\cos^2t+\sin^2t)}{sintcost}=\frac{sins}{sintcost}\]

myininaya (myininaya):

any questions?

OpenStudy (anonymous):

what happened to the sins in the last step? It should be a 2sins by the way I interpreted it\[sins(\cos^2t)+sins(\sin^2t)=2sins(\cos^2t+\sin^2t)\]

myininaya (myininaya):

where did you get 2?

myininaya (myininaya):

do you know how to factor?

myininaya (myininaya):

xy+xz does not equal 2x(y+z) xy+xz=x(y+z)

OpenStudy (anonymous):

oh, my goodness, I don't know why that wasn't obvious to me, but its very clear now. I get all of it. Really appreciate the help, Thanks!

myininaya (myininaya):

lol and you know that cos^2x+sin^2x=1 right? if you want to give me a medal thats cool lol

OpenStudy (anonymous):

yeah, the best identity there is!! haha. If you could point out how to do that, I will. Also, I don't know how to close this question as "Has a good answer"

myininaya (myininaya):

^^^^^^^^^ ||||||||||||||| the button is here

OpenStudy (anonymous):

i've clicked on the blue thing, your icon and the actual medal, nothing happens, I'm new here, like today new, if that matters.

myininaya (myininaya):

its cool no worries

myininaya (myininaya):

its above the word delete

OpenStudy (anonymous):

I see theres a good answer button in other open study windows I have, but not this window. I'll be asking help again sometime, and hopefully the button reappears by then lol.

myininaya (myininaya):

lol you did it! you get a medal for that

OpenStudy (anonymous):

there we go, just refreshed the page lol. Thanks!